OpenStudy (anonymous):

Events A and B are mutually exclusive. Also, P(A) = 0.15 and P(B) = 0.2. What is P(A or B)? A. 0.35 B. 0.03 C. 0.05 D. 0.20

OpenStudy (alyssa_xo):

P(A or B) = P(A) + P(B) (for mutually exclusive events)

OpenStudy (anonymous):

i'm lost O.o

OpenStudy (alyssa_xo):

the probability of event \(P(A or B)\) is the same thing as the probability of \(A\) happening and \(B\) happening. so \(P(A or B)=P(A)+P(B)\)

OpenStudy (alyssa_xo):

in our case: P(A) = 0.15 and P(B) = 0.2. P(A or B)=P(A)+P(B)

OpenStudy (anonymous):

so a ?

OpenStudy (alyssa_xo):

mhm

OpenStudy (anonymous):

ok :)

OpenStudy (alyssa_xo):

ooh a smiley she wants me guys